write the equation of the specified parabola:
1) vertex (0,0) and focus (0,3)
2) vertex (2,0) and focus (0,0)

Answer by venugopalramana(3286)

write the equation of the specified parabola:
STD. EQN. IF VERTEX AND FOCUS HAVE SAME X COORDIATE........IS
(X-H)^2=4A(Y-K)....WHERE (H,K) IS VERTEX AND (H,K+A) IS FOCUS
1) vertex (0,0) and focus (0,3)
STD. EQN. IF VERTEX AND FOCUS HAVE SAME X COORDIATE........IS
(X-H)^2=4A(Y-K)....WHERE (H,K) IS VERTEX AND (H,K+A) IS FOCUS
....H=0...K=0...K+A=3=0+A...A=3
(X-0)^2=4*3(Y-0)
X^2=12Y
2) vertex (2,0) and focus (0,0)
STD. EQN. IF VERTEX AND FOCUS HAVE SAME Y COORDIATE........IS
(Y-K)^2=4A(X-H)....WHERE (H,K) IS VERTEX AND (H+A,K) IS FOCUS
...H=2...K=0.....H+A=0....2+A=0....A=-2
(Y-0)^2=4*(-2)(X-2)
Y^2=-8(X-2)